# CraftRTL: High-quality Synthetic Data Generation for Verilog Code Models with Correct-by-Construction Non-Textual Representations and Targeted Code Repair

### v1

@@ 2.3 Variability on Pass Rates During Training
Figure 1: Agreement by pass rate between two subsequent checkpoints during training on SDG data for problems in VerilogEval-Human.
During our training, we observed significant variability in the model’s pass rate on specific benchmark problems across different checkpoints. We note such variance is different from training instability (Wortsman et al., 2023) as we observe a stable decrease in the training loss. This variability persists even in the later stages of training, despite using a low learning rate. We illustrate this variability in Figure 1. The scatter plot tracks the pass rate for each problem in VerilogEval-Human, with each point representing the pass rate for the same problem across two checkpoints. The size of each point indicates the number of problems with the same pass rates for the two model checkpoints. We further categorize the region into areas where the checkpoints agree on problem difficulty and areas where they do not.
@@
Figure 2: Generating non-textual problems and correct-by-construction solutions.
3.1 Ensuring Quality Through Correct-by-Construction
@@
We generate Verilog code problems and solutions that are correct-by-construction. Our focus is on creating problems and solutions for non-textual representations. The overview of the process for generating non-textual problems and correct-by-construction solutions is illustrated in Figure 2. Table 3 shows the quantity of our correct-by-construction generation data (referred to as CC). To prevent data contamination, we exclude entries that duplicate the data representations of benchmark problems.
@@
Algorithm 1 outlines the process for generating a Moore FSM with random transitions. State reachability is ensured by first constructing a tree. Legality for state transition is ensured by ensuring each node has an out-degree of $2^{w}$ with the input bit width of $w$. The result is an FSM where the transitions between states are randomly assigned but conform to the specified input bit width. The algorithm can be easily modified for a Mealy FSM by assigning the output to the edges rather than nodes.
@@ 3.2 Mitigating “Minor” Errors with targeted code repair
Figure 4: An overview of our approach in generating targeted code repair data.
@@
To systematically address the issue, we first created a comprehensive Error Report for benchmark problems using LLMs, targeting those with significant pass rate fluctuations across training checkpoints for models on SDG data. We prompt the LLM to examine the nature of the mistakes by comparing correct and erroneous code completions for each problem, categorizing the errors into functional types (such as logic, multiplexing, and arithmetic operations). Additionally, we prompt LLMs to identify mistakes in common error patterns, for example, frequent mistakes in arithmetic shifts, bit manipulation, and unintended latch creation. This detailed report not only categorizes the errors but also highlights areas where the model consistently underperforms.
@@
To ensure the reliability of the error report and the code repair dataset, we implemented a two-phase validation process with LLMs. In the first phase, we conducted a self-consistency check of the Error Report by having the language model attempt to fix the error code based on the report’s hints. This step verifies the accuracy of the report by confirming that the model can resolve the errors using the provided guidance. In the second phase, during the generation of the code repair dataset, we apply self-verification, including deduplication, syntax filtering, and benchmark decontamination. These measures ensure the dataset’s quality and uniqueness, preventing overlap with evaluation benchmarks.
@@ 4.1 Implementation Details
Table 4: We compare our models with various baseline models on VerilogEval (Liu et al., 2023b). We update the results from Zhao et al. (2024) with the latest foundational and frontier code models. The best results for each model type are highlighted in bold. We use red to denote the best model overall and blue to indicate the best results among models other than ours.
@@ 4.3 Results
Table 5: Evaluations on RTLLM v1.1 (Lu et al., 2024) using unbiased pass@k metrics. The best results for each model type are highlighted in bold. We use red to denote the best model overall and blue to indicate the best results among models other than ours. We re-evaluate all models (see Appendix A for details).
@@
Table 6: Ablation study on training data.
@@
Figure 5: Agreement by pass rate between two subsequent checkpoints during training on SDG-CC-Repair for problems in VerilogEval-Human.
Figure 5 displays the pass rates for two consecutive checkpoints of Starcoder2-SDG-CC-Repair on VerilogEval-Human problems, sampled with a temperature of 0.8. Compared to Figure 1, the updated model shows significant improvements by (1) moving previously unsolved problems into the solved category, including those with non-textual representations addressed by our correct-by-construction CC data, and (2) reducing the number of problems with large pass rate discrepancies, particularly where performance had degraded. The targeted repair data has effectively mitigated the model’s tendency to repeat common mistakes found in our Repair dataset, despite the noise inherent in synthetically generated SDG data.
@@
The quality control mechanisms integrated into the data generation pipeline are crucial for improving model performance, particularly in correcting minor errors through targeted code repair. To evaluate the impact of these quality controls, we conducted an ablation study in Table 7), where we systematically removed each component of the targeted code repair generation pipeline and assessed the resulting model performance. Specifically, we eliminated the self-consistency checks that validate whether the generated error report effectively guides the LLMs in correcting mistakes. Additionally, we tested the removal of the error report entirely, substituting it with random errors injected into the open-source code by the LLMs. The benchmark results indicate a significant performance drop when these validation processes are excluded. These findings highlight the essential role of both the self-consistency checks and the targeted error report in improving the model’s ability to correct errors.
@@ 6 Discussions
In this work, we refer to synthetic data generation as methods of using large language models (LLMs) in data generation. While our approach—ensuring correctness through correct-by-construction—could also be considered “synthetic” and resembles methods explored in works like AlphaGeometry (Trinh et al., 2024), our problems are much simpler and on a smaller scale. Our observations about the variability of models on specific problems align with the findings of  Meta AI (2024), where “the model knows how to produce the right answer, but it does not know how to select it.” Instead of striving for absolute data correctness, preference learning (Rafailov et al., 2024; Ethayarajh et al., 2024) or reinforcement learning (Bai et al., 2022; Le et al., 2022), we generate targeted repair data by analyzing errors and re-create such scenarios by injecting similar errors into open-source code, somewhat analogous to how humans consolidate memories during sleep by integrating new information with past experiences (Walker & Stickgold, 2004; Stickgold, 2005).
@@ A.1 Our Models
We present our models’ results on Verilog benchmarks tested with temperatures 0.2 and 0.8. We ablate across different data blends, with SDG indicating using LLM synthetic generated data in Section 2.1, CC indicating correct-by-construction data targeting non-textual representations in Section 3.1, and Repair representing our targed code repair dataset in Section 3.2.
@@ A.8 Diversity of Generated Code
Appendix B Examples of Targeted Code Repair Data
@@
Appendix C Examples of Correct-by-Construction Data for Non-Textual Representations
C.1 Karnaugh Maps and Truth Tables
@@
C.2 State Transition Graphs and Tables
@@ C.3 Waveforms
Appendix D Prompt Templates
D.1 Synthetic Data Generation
D.1.1 Self-Instruct
@@ D.1.4 Non-textual Representations
D.1.5 Prompts for Sampling Solutions with LLM Generated Problems

### v2

@@ 2.2 Challenges with Non-Textual Representations
(a) Starcoder2-15B on SDG.
(b) Starcoder2-15B on SDG-CC-Repair.
Figure 1: Our methods reduce pass rate variability during training: SDG (left) shows high volatility with significant degradation on many problems, while SDG-CC-Repair (right) stabilizes learning outcomes on solvable problems (details in Section A.10).
@@ 2.3 Variability on Pass Rates During Training
During our training, we observed significant variability in the model’s pass rate on specific benchmark problems across different checkpoints. We note such variance is different from training instability (Wortsman et al., 2023) as we observe a stable decrease in the training loss. This variability persists even in the later stages of training, despite using a low learning rate. We illustrate this variability in Figure 1(a). The scatter plot tracks the pass rate for each problem in VerilogEval-Human, with each point representing the pass rate for the same problem across two checkpoints. The size of each point indicates the number of problems with the same pass rates for the two model checkpoints. We further categorize the region into areas where the checkpoints agree on problem difficulty and areas where they do not.
@@
3.1 Ensuring Quality Through Correct-by-Construction
We generate Verilog code problems and solutions that are correct-by-construction. Our focus is on creating problems and solutions for non-textual representations. Table 3 shows the quantity of our correct-by-construction generation data (referred to as CC). To prevent data contamination, we exclude entries that duplicate the data representations of benchmark problems.
@@ 3.1 Ensuring Quality Through Correct-by-Construction
Algorithm 1 outlines the process for generating a Moore FSM with random transitions. State reachability is ensured by first constructing a tree. Legality for state transition is ensured by ensuring each node has an out-degree of $2^{w}$ with the input bit width of $w$. The result is an FSM where transitions between states are randomly assigned but conform to the specified input bit width. The algorithm can be easily modified for a Mealy FSM by assigning the output to the edges rather than nodes.
@@ 3.2 Mitigating “Minor” Errors with targeted code repair
Figure 3: Overview of our approach for generating targeted code repair data: (1) prompting the LLM to generate detailed error reports from correct and erroneous code, (2) validating error report quality by ensuring the LLM can debug the errors based on the report, and (3) leveraging the LLM to inject similar errors into open-source code, creating a diverse training dataset.
@@
To systematically address the issue, we first created a comprehensive Error Report for benchmark problems using LLMs, targeting those with significant pass rate fluctuations across training checkpoints for models on SDG data. We prompt the LLM to examine the nature of the mistakes by comparing correct and erroneous code completions for each problem, categorizing the errors into common error types (details in Section A.9). This detailed report not only categorizes the errors but also highlights areas where the model consistently underperforms.
@@
To ensure the reliability of the error report and the code repair dataset, we implemented a two-phase validation process with LLMs. In the first phase, we conducted a self-consistency check of the Error Report by having the language model attempt to the fix error code based on the report’s hints. This step verifies the accuracy of the report by confirming that the model can resolve the errors using the provided guidance, whereas directly prompting the LLM without detailed error reports could resolve only 13% of the errors. In the second phase, during the generation of the code repair dataset, we apply self-verification, including deduplication, syntax filtering, and benchmark decontamination. These measures ensure the dataset’s quality and uniqueness, preventing overlap with evaluation benchmarks.
@@ 4.1 Implementation Details
Table 4: We compare our models with various baseline models on VerilogEval (Liu et al., 2023b). We update the results from Zhao et al. (2024) with the latest foundational and frontier code models. The best results are highlighted in bold.
@@ 4.3 Results
Table 5: Evaluations on RTLLM v1.1 (Lu et al., 2024) using unbiased pass@k metrics. The best results are highlighted in bold. We re-evaluate all models (see Appendix A for details).
@@
Table 6: Ablation study on training data. Data quantity indicated in parentheses.
@@
Figure 1(b) displays the pass rates for two consecutive checkpoints of Starcoder2-SDG-CC-Repair on VerilogEval-Human problems, sampled with a temperature of 0.8. Compared to Figure 1(a), the updated model shows significant improvements by (1) moving previously unsolved problems into the solved category, including those with non-textual representations addressed by our correct-by-construction CC data, and (2) reducing the number of problems with large pass rate discrepancies, particularly where performance had degraded. The targeted repair data has effectively mitigated the model’s tendency to repeat common mistakes found in our Repair dataset, despite the noise inherent in synthetically generated SDG data.
@@
0055551010101015151515202020202525252530303030000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.0500.9560.9650.9700.9861.001.001.000.0000.7600.7990.8460.9310.9620.9941.000.6450.6880.7100.7220.7310.7510.7610.1030.5510.5790.5930.6390.6580.6760.664Total Number of Targeted Data Samples (k)pass@1KMapFSMWaveformVerilogEval-NonText
@@
The quality control mechanisms integrated into the data generation pipeline are crucial for improving model performance, particularly in correcting minor errors through targeted code repair. To evaluate the impact of these quality controls, we conducted an ablation study in Table 7, where we systematically removed each component of the targeted code repair generation pipeline and assessed the resulting model performance. Specifically, we eliminated the self-consistency checks that validate whether the generated error report effectively guides the LLMs in correcting mistakes. Additionally, we tested the removal of the error report entirely, substituting it with random errors injected into the open-source code by the LLMs. The benchmark results indicate a significant performance drop when these validation processes are excluded. These findings highlight the essential role of both the self-consistency checks and the targeted error report in improving the model’s ability to correct errors.
@@ 6 Discussions
In this work, we refer to synthetic data generation as methods of using large language models (LLMs) in data generation. While our approach—ensuring correctness through correct-by-construction—could also be considered “synthetic” and resembles methods explored in works like AlphaGeometry (Trinh et al., 2024), our problems are much simpler and on a smaller scale. Our observations about the variability of models on specific problems align with the findings of  Meta AI (2024), where “the model knows how to produce the right answer, but it does not know how to select it.” Instead of striving for absolute data correctness, preference learning (Rafailov et al., 2024; Ethayarajh et al., 2024) or reinforcement learning (Bai et al., 2022; Le et al., 2022), we generate targeted repair data by analyzing errors and re-create such scenarios by injecting similar errors into open-source code, somewhat analogous to how humans consolidate memories during sleep by integrating new information with past experiences (Walker & Stickgold, 2004; Stickgold, 2005). Further discussions on the generalizability and broader impact of our work are provided in Appendix B.
@@ 7 Conclusion
Reproducibility Statement
We provide the following details: evaluation benchmarks in Section A.3, examples of the process for generating targeted code repair data in Appendix C, and data examples from correct-by-construction targeting non-textual representations in Appendix D. Additionally, we include prompt templates used for data generation in Appendix E. Our data generation pipeline is available: https://github.com/NVlabs/CraftRTL.
@@ A.1 Our Models
We present our models’ results on Verilog benchmarks tested with temperatures 0.2 and 0.8. We ablate across different data blends, with SDG indicating using LLM synthetic generated data in Section 2.1, CC indicating correct-by-construction data targeting non-textual representations in Section 3.1, and Repair representing our targeted code repair dataset in Section 3.2.
@@ A.8 Diversity of Generated Code
A.9 Error Types of LLM generated Error reports
Table 15: Error types of LLM generated error reports.
Error Type #Errors One-line Description
Vector Concatenation 15.3% Errors during vector concatenation or bit slicing.
Incorrect Initialization 13.1% Missing or faulty initialization of registers or signals.
Boolean Logic Flaws 12.4% Logical inconsistencies or errors in combinational logic expressions.
Shift Operation Faults 10.2% Misaligned or unintended behavior during shift operations.
Timing Violations 10.2% Errors where signal propagation violates timing requirements.
KMap Misinterpretation 8.8% Incorrect derivation of Boolean expressions from Karnaugh maps.
Latch Hazards 6.5% Unintended latches caused by missing or faulty conditions.
Bit Manipulation Bugs 7.3% Errors in operations like masking, flipping, or extracting specific bits.
Casez Priority Conflicts 4.4% Ambiguities or conflicts in casez or case statements.
Nested Loop Design Flaws 3.7% Incorrect or inefficient nested loop designs.
Others 8.1% Miscellaneous errors not covered above.
Table 15 shows the distribution of common error types in LLM-generated error reports, along with brief one-line descriptions. Most of these “minor” errors occur in solvable problems and stem from hardware-specific concepts (e.g., shift operations, timing violations) and Verilog related issues uncommon in software languages (e.g., latch hazards, casez priority conflicts). When generating targeted repair training data, we randomly sample detailed error reports and open-source code snippets, ensuring the error type distribution in training aligns with their natural occurrences.
A.10 Details on Figure 1
Table 16: Checkpoints of Figure 1.
Model checkpoint1 checkpoint2
Steps Epoch Steps Epoch
SDG 256 0.82 320 1.0
SDG-CC-Repair 386 0.86 448 1.0
In Section 2.3 we discussed our findings on training variability in learning outcomes for specific benchmark problems. To analyze this, we saved checkpoints every 64 gradient steps during training and tracked the pass rates of specific benchmarks. Our training process is limited to a single epoch, as further training was found to be not helpful. We classify problems with pass rates exceeding 67% as solvable, and those below 33% as unsolvable. For the visualizations in Figure 1 we selected the final two saved checkpoints, detailed in Table 16. The ideal outcome is not merely reduced variability but also less degradations and improved accuracy: specifically, most problems in checkpoint2 should show higher pass rates than checkpoint1, assuming that training on additional data enhances model performance. However, as shown in Figure 1(a) training on SDG data results in a significant degradation of pass rates for many problems between checkpoint1 and checkpoint2. In contrast Figure 1(b) demonstrates reduced degradation and improvement in more problems. We further elaborate such findings in Table 17, where we display pass rates for selected benchmark problems with high volatility from VerilogEval-Human throughout the training progression.
Table 17: We displays pass rates for selected benchmark problems from VerilogEval-Human throughout the training progression. Each entry shows the pass rate for SDG-CC-Repair (SDG), with SDG results in parentheses.
Problem Step 64 Step 128 Step 256 Step 320 Step 386 Step 448
m2014_q4h 1.0 (1.0) 1.0 (0.9) 1.0 (0.967) 1.0 (0.875) 1.0 (-) 1.0 (-)
always_nolatches 1.0 (0.867) 1.0 (0.9) 1.0 (0.6) 1.0 (0.833) 1.0 (-) 1.0 (-)
vectorr 1.0 (0.633) 1.0 (0.925) 1.0 (0.467) 0.95 (0.925) 1.0 (-) 1.0 (-)
fsm2s 1.0 (0.8) 1.0 (0.8334) 0.8 (0.775) 1.0 (0.967) 1.0 (-) 1.0 (-)
fsm3comb 1.0 (0.0) 0.95 (1.0) 0.5 (0.533) 1.0 (0.233) 1.0 (-) 1.0 (-)
We believe such volatility primarily is due to noise in SDG data where we can not verify solution correctness. Because of the difficulties of verifying coding solutions in hardware descriptive languages, we instead generate targeted repair data for LLMs to learn to mitigate common errors which have shown to generalize to writing correct code during completion. To the best of our knowledge, we are the first work to describe such findings and provide an effective solution.
Appendix B Further Discussions and Broader Impacts
In this section, we provide further discussions to address concerns regarding the novelty, generalizability, and significance of our proposed methods. We offer clarifications to highlight the relevance and broader impact of our work, underscoring its value to the broad research community.
B.1 Generalizability of Correct-by-Construction Data Generation
Our approach to curating correct-by-construction data is largely inspired by Trinh et al. (2024), who introduced a mathematically rigorous method utilizing symbolic deduction engines to construct synthetic training data, significantly improving LLM capabilities in solving Olympiad geometry problems. Similarly, our method ensures the correctness of problems and solutions through a custom-designed data generation pipeline, leveraging custom-designed solvers to generate accurate solutions to their corresponding problems. In contrast to methods distilling LLM responses like Self-Instruct (Wang et al., 2022), our correct-by-construction approach ensures data quality and solution accuracy without relying on strong LLM performance on downstream tasks. We hope that our mathematically rigorous approach to generating synthetic data can further inspire future work on improving LLMs general capabilities in areas such as math, coding, and symbolic reasoning. Moreover, we recognize that adapting these methods to other domains may require human tuning to identify the best data generation method, and we note that automating this process for scalability could be a promising future research direction.
B.2 Novelty and Generalizability of Targeted Code Repair
Our analysis show that LLMs frequently make “minor” errors in Verilog coding, often correctable within few lines of code. We attribute this primarily to the LLMs’ insufficient training in comprehending problem descriptions and instructions alongside their correct solutions. Prior research has tackled this challenge by improving data quality. For instance, Chen et al. (2022) filters incorrect code using tests generated by LLMs, while Zhang et al. (2024c) creates preference learning datasets by ranking code through self-validation. Lei et al. (2024) focus on generating fine-tuning data through code completion, test validation, and debugging with LLM agents, while Le et al. (2022) trained reward models based on compilation and unit test outcomes to enhance LLM performance via reinforcement learning. However, low-resource languages face additional obstacles due to limited data availability, making it particularly difficult to synthesize unit tests directly in these languages. To address this issue, Cassano et al. (2024) introduced lightweight compilers to translate test cases from source to target languages.
Verilog coding encounters challenges typical of low-resource languages, compounded additional domain-specific challenges as a hardware description language rather than a conventional programming language. Its unique characteristics pose significant barriers to knowledge transfer from high-resource languages, as highlighted in studies on execution performance in parallel programming (Nichols et al., 2024) and high-performance computing extensions (TehraniJamsaz et al., 2024). To address these challenges, we propose a novel pipeline for generating targeted code repair data. While automatic code repair has been extensively studied, most existing methods focus on widely-used programming languages (Xia et al., 2023), relying on data of buggy code and fixes from open-source repositories (Tufano et al., 2019; Just et al., 2014). In contrast, our pipeline utilizes a small set of well-curated benchmarks and testbench to automate the generation of error reports, quality assurance, and augmentation of training datasets by injecting similar errors into open-source code. Our results highlight the effectiveness of this approach, which is language agnostic and can be adapted to other low-resource and domain-specific programming languages.
B.3 Significance of Non-Textual Data Representations in Hardware Design
In this work, we emphasize the significance of non-textual data representations, specifically Karnaugh maps, state-transition diagrams, and waveforms, for accurately capturing hardware functionality. These representations are widely utilized by hardware designers to mitigate the ambiguity and verbosity inherent in natural language descriptions. While they may be specific to hardware design, they are not limited to Verilog and can be applied to various domain-specific languages (DSLs) for hardware design. This is supported by Batten et al. (2024), who leveraged similar non-textual representations from VerilogEval-Human to evaluate the performance of LLMs on several Python-embedded hardware design DSLs.
In this study, we focus exclusively on limited representations, which constitute a significant portion (45 problems, approximately 30%) of all problems in the VerilogEval-Human benchmark (details in Section A.4). We exclude other types of non-textual representations due to the lack of a suitable benchmark for evaluating LLMs in Verilog coding. Chang et al. (2024) emphasize the importance of non-textual representations, particularly visual representations, in describing hardware designs. While their work targets visual-language models and is therefore beyond the scope of this study, we recognize that similar methodologies of our work such as correct-by-construction methods could be employed to generate training data for visual representations, such as circuit schematics, data flow diagrams, and state transition graphs.
Appendix C Examples of Targeted Code Repair Data
@@
Appendix D Examples of Correct-by-Construction Data for Non-Textual Representations
D.1 Karnaugh Maps and Truth Tables
@@
D.2 State Transition Graphs and Tables
@@ D.3 Waveforms
Appendix E Prompt Templates
E.1 Synthetic Data Generation
E.1.1 Self-Instruct
@@ E.1.4 Non-textual Representations
E.1.5 Prompts for Sampling Solutions with LLM Generated Problems